# Canonical formalism in the Schwarzschild spacetime

### v2

@@ I.3.1 Geometric generalities
To apply the general formalism reviewed in Sec. I.1 to the case of the Schwarzschild spacetime, we make a choice of coordinates $x^{\alpha}$ covering that spacetime. We choose the Schwarzschild-Droste coordinates $x^{\alpha}=(t,r,\theta,\phi)$, in which the contravariant components of the metric read (cf. Ref. [85])
@@ I.3.2 Invariants from Killing vectors
Next we look at the symmetries of the Schwarzschild spacetime, which are in correspondence with four Killing vector fields. Given in the natural vector basis $(\partial_{\alpha})^{a}$ associated with $x^{\alpha}=(t,r,\theta,\phi)$, they read (cf. Ref. [87])
@@
It follows from the definitions in Eq. (41) and the Hamiltonian in Eq. (35) that $E$, $L_{z}$ and $L^{2}$ are all first integrals of the system on the $12$D phase space $\mathcal{N}$ (i.e., functions $\mathcal{I}:\mathcal{N}\rightarrow\mathbb{R}$ such that $\{\mathcal{I},H_{\mathcal{N}}\}=0$). While this is rather trivial for $E$ and $L_{z}$ because $H_{\mathcal{N}}$ in Eq. (35) is independent of $t$ and $\phi$ (i.e. the latter are “cyclic” coordinates), the fact that $L^{2}$ is a first integral follows from an explicit (but easy) calculation of the Poisson bracket $\{L^{2},H_{\mathcal{N}}\}=0$ with Eq. (26).
@@ I.3.3 Invariants from Killing-Yano tensors
In the Schwarzschild spacetime, there exists a rank-2 Killing-Yano tensor $Y^{ab}$. In the natural basis, it is given by (cf. Refs. [89, 90])
@@
$\{\mathfrak{K},H_{\mathcal{N}}\}=-\sqrt{f}\,{C}^{0}\neq 0\,,\quad\text{and}\quad\{\mathfrak{Q},H_{\mathcal{N}}\}=0\,.$ (45)
@@
Our first observation in Eq. (45) is that $\mathfrak{K}$ is not an invariant of the $12$D phase space $\mathcal{N}$ because $C^{0}$ on the right-hand side does not vanish there. In fact, $\mathfrak{K}$ will only be invariant on the $10$D physical phase space $\mathcal{P}$, where the TD SSC is satisfied globally. This should be contrasted with the four Killing-vector invariants in Eq. (38) that are first integrals on $\mathcal{N}$ and on $\mathcal{P}$. Our second observation in Eq. (45) is that $\mathfrak{Q}$ is a first integral on $\mathcal{N}$. Although remarkable and unexpected (since $\mathfrak{Q}$ is invariant only under the TD SSC. See Ref. [80] for further more details), this is a simple consequence of the equality
@@
which follows from comparing Eqs. (41c) and (44b). This means that the Killing-tensor invariant $\mathfrak{Q}$ corresponds to the Killing-vector invariant $L^{2}$. Why this coincidence occurs is a deep question, for which we currently have no convincing answer. [^8] Remarkably, the fact that $\mathfrak{Q}$ is a conserved quantity on $\mathcal{N}$ is also true in Kerr, even though there one does not have Eq. (46) to justify it (there is no conserved angular-momentum vector in Kerr. This is unexpected, since Rüdiger invariants were built in [78, 79, 80] while enforcing the TD SSC. Yet, the invariance of $\mathfrak{Q}$ does not seem to need the TD SSC, in contrast with $\mathfrak{K}$. It is all the more surprising that Eq. (46) comes from the fact that any spin dependent term other than that proportional $S^{1}$ are exactly cancelled out in Eq. (16b), again without any constraints in Eq. (36). The only worthwhile comment we can make is that the correspondence in Eq. (46) is probably reminiscent of another well-known fact in the non-spinning limit; in the case of Schwarzschild spacetime, the Carter constant $K:=K_{ab}p^{a}p^{b}$ (cf. Ref. [45]), which is identical to the “non-spinning part” of $\mathfrak{Q}$, also agrees with the total orbital angular momentum $L^{2}=p^{2}_{\theta}+\csc^{2}\theta\,p_{\phi}^{2}$ [cf. App. B].
We conclude that the Killing-Yano tensor in Schwarzschild spacetime gives rise to only one non-trivial Rüdiger invariant $\mathfrak{K}$ exclusively at the level of $\mathcal{P}$, not $\mathcal{N}$. The other invariant $\mathfrak{Q}$ is conserved on $\mathcal{N}$ but it is equivalent to the Killing-vector invariant $L^{2}$, already introduced in the preceding subsection.
@@ I.3.4 Angular momentum and spin configurations
It is perhaps helpful to note that $\mathfrak{K}$ and $\mathfrak{Q}$ in the physical phase space $\mathcal{P}$ can be given an alternative representation in terms of a covariant notion of angular-momentum four vector (cf. Sec. IV.C.3 in Paper I), namely
@@
constructed from the Killing-Yano tensor $Y^{ab}$ (cf. Refs. [91, 92]). The vector ${\mathcal{L}}^{a}$ has the dimension of angular momentum $[\text{length}]^{+2}$, and its explicit expression in the Schwarzschild spacetime is
@@
The expressions for $\mathfrak{K}$ and $\mathfrak{Q}$ in Eq. (49) sometimes offer a practical advantage over those in Eq. (44) because their interpretation is intuitively clear. The form of ${\cal L}^{2}$ bears a formal resemblance to the Carter constant of the non-spinning body in the linearized Kerr spacetime: ${\mathcal{L}}^{2}=(C+L_{z}^{2})-2EaL_{z}+O(a^{2})$, where $a$ is the Kerr parameter and $C:=K-(L_{z}-aE)^{2}$ is an alternative form of the Carter constant [cf. Refs [94, 95, 96]]. As the form suggests, $\mathfrak{K}$ represents the spin-orbit coupling between $S_{a}^{\text{\tiny{TD}}}$ and ${\mathcal{L}}^{a}$. [^9] This may be viewed as a relativistic extension of the conserved spin-orbit coupling identified in the post-Newtonian formalism (cf. Refs. [97, 98]). In particular, the correspondence (49) motivates us to introduce special configurations for the spin $S^{a}_{\text{\tiny{TD}}}$, depending on whether it is perpendicular to or aligned with the angular momentum ${\mathcal{L}}^{a}$. Taking Eqs. (13) and (49) into account and that $L^{2}=\mathfrak{Q}={\cal L}^{2}+O(\epsilon)$ with Eq. (46), we then define the following spin configurations:
@@ II Hamiltonian system on $\mathcal{N}$ and Andoyer variables
Andoyer variables, named after the French astronomer Henri Andoyer (1862-1929), are a set of canonical coordinates used in the Hamiltonian mechanics of Newtonian rigid body motion: see, e.g., Refs. [99, 100]. They are a generalization of the more well-known Hill and Delaunay variables used in the reduction of the Newtonian two-body problem (cf. Refs. [101, 62]). The introduction of Andoyer-type variables in general relativity for spinning dynamics presented here is new, and similar to the definition of orbital elements defined in the geodesic case in Sec. II.1, although more geometry is involved. Quite remarkably, this “classical” change of coordinates can also be applied to the present relativistic context. It is, in essence, the only means by which we are practically able to transition to the physical phase space $\mathcal{P}$.
@@ II.1 Schwarzschild geodesics
The motion of a non-spinning test particle in a given background spacetime is described by geodesics. These geodesics have a well-known Hamiltonian formulation (cf. Refs. [46, 53]) on a $8$D phase space $\mathcal{G}$ (cf. FIG. 1), spanned by the $4$ spacetime coordinates $x^{\alpha}$ and their $4$ conjugated momenta $p_{\alpha}$, corresponding to the particle’s momentum $p_{a}$ components. The Hamiltonian generating the geodesic equations is simply $H_{\mathcal{G}}:=\tfrac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}$. For the case of the Schwarzschild spacetime, this reads
@@
where $(t,r,\theta,\phi)$ are the Schwarzschild-Droste coordinates and $(p_{t},p_{r},p_{\theta},p_{\phi})$ are the components of the $1$-form $p_{a}$ in the natural basis $(\partial_{\alpha})^{a}$. Just like the Kepler problem in Newtonian gravity (of which the Hamiltonian is given by the leading order in a $M/r\ll 1$ expansion of Eq.(52)), Schwarzschild geodesics are always spatially confined to a $2$-dimensional plane, reflecting the spherical symmetry. In terms of Hamiltonian mechanics, this geometric feature is nothing but a $1:1$ resonance between the $\theta$ and $\phi$ motion; this fact is not so obvious by direct inspection of Eq. (52). Because the confinement is rarely presented as a resonance in relativistic mechanics, we present a detailed discussion of it in App. B.2.
@@
where $\nu$ is the angle defining the line of nodes, where the orbital plane intersects the equatorial plane, and $\ell$ is the true anomaly, which specifies the location of the particle within the orbital plane; see Fig. 2. With the help of some Euclidean trigonometry, as detailed in App. B, we find the following formulae
@@ II.2 Construction of relativistic Andoyer variables
Having reviewed the non-spinning (geodesic) case in the previous section, we now generalize the transformation in Eq. (53) to the case of a spinning body (consistently at linear order in spin). In addition to $\theta,\phi$, such transformation must now involve the spin angle $\sigma$, since the angular momentum $\pi_{\ell}$ depends on its conjugated momentum $\pi_{\sigma}$, cf. Eq. (41c). Therefore, our goal is to define three new angles $(\nu,\ell,s)$ such that their conjugated momenta are related to the angular momenta of the system by $\pi_{\nu}=\pi_{\phi},\,\pi_{s}=\pi_{\sigma}$ and $\pi_{\ell}=L$, where $L$ is the norm of the angular momentum $3$-vector in Eq. (41c). Other choices are possible, but this is probably the most simple one, and turns out to be close to a well-known transformation in classical mechanics.
@@
First, we define the set of Andoyer angles $(\nu,\ell,s)$, where $\nu$ and $\ell$ are the same angles as in the geodesic construction (cf. Sec. II.1), while $s$ belongs to the Body plane. The Andoyer angles can be related to the old ones $(\theta,\phi,\sigma)$ using spherical trigonometry identities (obtained in the spherical triangle formed by the three pairwise intersections of each plane, as in Fig. 3). These read  [^11] Compared to Ref. [100], we have shifted the angle $\ell$ by $\pi/2$ so that it coincides with its usual construction. This does not break the symplecticity because we have $\{\ell,F\}=\{\ell+\pi/2,F\}$ for any function $F$ of the phase space.
@@
The whole set of relations in Eqs. (59), (60) and (61) defines a canonical change of coordinates that generalizes the geodesic one in Eq. (53), namely
$(\theta,\pi_{\theta},\phi,\pi_{\phi},\sigma,\pi_{\sigma})\mapsto(\ell,\pi_{\ell},\nu,\pi_{\nu},s,\pi_{s})\,.$ (62)
To prove this, it suffices to either compute the symplectic form just like Andoyer did in classical mechanics (see, e.g., Ref. [100]), or to use brute force and check that all Poisson brackets remain canonical. It is worth noting that some formulae, like Eqs. (59) or Eq. (60), only define the new Andoyer variables up to signs. These signs are all fixed by asking that the transformation be canonical. [^12] For example, only the square of $\pi_{\theta}$ can be extracted from Eq. (60) After imposing $\{\theta,\pi_{\theta}\}=1$, one readily finds $\pi_{\theta}=-\pi_{\ell}\sin\iota\cos(\nu-\phi)$. Similar features happen for other variables.
@@ II.3 Interpretation of Andoyer variables
The Andoyer variables of (non-relativistic) classical mechanics have to do with the evolution of a free rigid body (cf. Refs. [102, 99, 100].) In our setup, they do not have quite the same Euclidean interpretation, and it is rather the mathematical formulae that they satisfy that we exploit. In this section we get a grasp at what our Andoyer-type variables represent for the relativistic system at hand.
@@ II.3.2 Planar (and equatorial) orbits
Let us now examine the $\pi_{s}=0$ limit, bearing in mind that this does not set all the spin terms to zero, cf. Eq. (28d). In this limit Eqs. (59) become  [^13] Just like the inclination angle $\iota$, we assume (without loss of generality) that $\delta\in[0,\pi]$, so that $\pi_{s}=0$ is equivalent to $\cos\delta=0$ (cf. Eq. (61)), implying both $\delta=\pi/2$ and $\sin\delta=1$.
@@
$\sin(\sigma-s)\sin\theta$ $=\sin\iota\cos\ell\,,$ (64d)
$\cos(\sigma-s)\sin\theta$ $=\cos\iota\,.$ (64e)
In the first three equations, we identify the geodesic formulae in Eqs. (54), for which the motion was planar. Indeed, if $\pi_{s}=0=\pi_{\sigma}$, then it follows from Eq. (58) that $\vec{L}$ is orthogonal to $\vec{e}_{r}$ at all times: the motion is confined into the invariant Andoyer plane (orthogonal to $\vec{L}$). Consequently, it appears that the variable $\pi_{s}=\pi_{\sigma}$ controls the non-planarity of the orbit. In particular, the angle between the position vector $\vec{e}_{r}$ and the Andoyer plane is $\delta-\pi/2$.
The angle $s$ is measured within the Body plane, and locates the position of the spin vector $\vec{S}$ (cf Eq. (19)) with respect to the intersection of the Body and Andoyer plane. The angle $s$ coincides with the spin angle $\sigma$ only for equatorial orbits, where one has $(\theta,\phi,\iota,\sigma)=(\pi/2,\ell+\nu,0,0)$. Therefore, $s$ is nothing but the generalization of the angle $\sigma$ parametrizing the spin vector $\vec{S}$ in the spherical triad, and we refer to Sec. III.B in Paper I for its physical interpretation.
Lastly, the dynamics will reveal that if $\pi_{s}=0$, then $s=0$ as well, as seen in Sec. II.5.
Although the orbit of the spinning body is generally non-planar, the dynamics still exhibit an invariant plane, which is orthogonal to the constant angular momentum vector $\vec{L}$. This plane is characterized by a line of nodes $\nu$ and an inclination $\iota$ (cf. Eq. (61)). We stress, again, that the motion of the spinning body is not confined to this plane, but stays $O(\epsilon)$-close to it.

### v3

@@ I.3.1 Geometric generalities
To apply the general formalism reviewed in Sec. I.1 to the case of the Schwarzschild spacetime, we make a choice of coordinates $x^{\alpha}$ covering that spacetime. We choose the Schwarzschild-Droste coordinates $x^{\alpha}=(t,r,\theta,\phi)$, in which the contravariant components of the metric read [cf. Ref. [116]]
@@ I.3.2 Invariants from Killing vectors
Next we look at the symmetries of the Schwarzschild spacetime, which are in correspondence with four Killing vector fields. Given in the natural vector basis $(\partial_{\alpha})^{a}$ associated with $x^{\alpha}=(t,r,\theta,\phi)$, they read [cf. Ref. [118]]
@@
It follows from the definitions in Eq. (42) and the Hamiltonian in Eq. (36) that $E$, $L_{z}$ and $L^{2}$ are all first integrals of the system on the $12$D phase space $\mathcal{N}$ (i.e., functions $\mathcal{I}:\mathcal{N}\rightarrow\mathbb{R}$ such that $\{\mathcal{I},H_{\mathcal{N}}\}=0$). While this is rather trivial for $E$ and $L_{z}$ because $H_{\mathcal{N}}$ in Eq. (36) is independent of $t$ and $\phi$ (i.e. the latter are “cyclic” coordinates), the fact that $L^{2}$ is a first integral follows from an explicit (but easy) calculation of the Poisson bracket $\{L^{2},H_{\mathcal{N}}\}=O(\epsilon^{2})$ with Eq. (27).
@@ I.3.3 Invariants from Killing-Yano tensors
In the Schwarzschild spacetime, there exists a rank-2 Killing-Yano tensor $Y^{ab}$. In the natural basis, it is given by [cf. Refs. [120, 121]]
@@
$\{\mathfrak{K},H_{\mathcal{N}}\}=-\sqrt{f}\,{C}^{0}+O(\epsilon^{2})\neq 0\,,\quad\text{and}\quad\{\mathfrak{Q},H_{\mathcal{N}}\}=O(\epsilon^{2})\,.$ (46)
@@
Our first observation in Eq. (46) is that $\mathfrak{K}$ is not an invariant of the $12$D phase space $\mathcal{N}$ because $C^{0}$ on the right-hand side does not vanish there. In fact, $\mathfrak{K}$ will only be invariant on the $10$D physical phase space $\mathcal{P}$, where the TD SSC is satisfied globally. This should be contrasted with the four Killing-vector invariants in Eq. (39) that are first integrals on $\mathcal{N}$ and on $\mathcal{P}$. Our second observation in Eq. (46) is that $\mathfrak{Q}$ is a first integral on $\mathcal{N}$999 Although unexpected (since $\mathfrak{Q}$ was built to be invariant only under the TD SSC [109, 110, 111], we observe that this feature is but a consequence of the equality
@@
which follows from comparing Eqs. (42c) and (45b). This means that the Killing-tensor invariant $\mathfrak{Q}$ corresponds to the Killing-vector invariant $L^{2}$. The latter being an invariant irrespective of any SSC, the equality (47) explains the result in (46).
Although one might be inclined to posit that the Schwarzschild spherical symmetry can be used to explain Eq.  (47) and thus the result in Eq.  (46), it is important to note that the latter is also true for the Kerr spacetime [cf. paper I. [1]], where one loses spherical symmetry and one does not have Eq. (47). We shall leave any justification of this unexpected fact for the time being.
The only worthwhile comment we can make is that the correspondence in Eq. (47) is probably reminiscent of another well-known fact in the non-spinning limit; in the case of Schwarzschild spacetime, the Carter constant $K:=K_{ab}p^{a}p^{b}$ [cf. Ref. [67]], which is identical to the “non-spinning part” of $\mathfrak{Q}$, also agrees with the total orbital angular momentum $L^{2}=p^{2}_{\theta}+\csc^{2}\theta\,p_{\phi}^{2}$ [cf. App. A].
Regardless, we conclude that the Killing-Yano tensor in Schwarzschild spacetime gives rise to only one non-trivial Rüdiger invariant $\mathfrak{K}$ exclusively at the level of $\mathcal{P}$, not $\mathcal{N}$. The other invariant $\mathfrak{Q}$ is conserved on $\mathcal{N}$ but it is equivalent to the Killing-vector invariant $L^{2}$, already introduced in the preceding subsection.
@@ I.3.4 Angular momentum and spin configurations
It is perhaps helpful to note that $\mathfrak{K}$ and $\mathfrak{Q}$ in the physical phase space $\mathcal{P}$ can be given an alternative representation in terms of a covariant notion of angular-momentum four vector [cf. Sec. IV.C.3 in Paper I], namely
@@
constructed from the Killing-Yano tensor $Y^{ab}$ [cf. Refs. [122, 123]]. The vector ${\mathcal{L}}^{a}$ has the dimension of angular momentum $[\text{length}]^{+2}$, and its explicit expression in the Schwarzschild spacetime is
@@
The expressions for $\mathfrak{K}$ and $\mathfrak{Q}$ in Eq. (50) sometimes offer a practical advantage over those in Eq. (45) because their interpretation is intuitively clear. The form of ${\cal L}^{2}$ bears a formal resemblance to the Carter constant of the non-spinning body in the linearized Kerr spacetime: ${\mathcal{L}}^{2}=(C+L_{z}^{2})-2EaL_{z}+O(a^{2})$, where $a$ is the Kerr parameter and $C:=K-(L_{z}-aE)^{2}$ is an alternative form of the Carter constant [cf. Refs. [125, 126, 127]]. As the form suggests, $\mathfrak{K}$ represents the spin-orbit coupling between $S_{a}^{\text{\tiny{TD}}}$ and ${\mathcal{L}}^{a}$. [^10] This may be viewed as a relativistic extension of the conserved spin-orbit coupling identified in the post-Newtonian formalism [cf. Refs. [128, 129]]. In particular, the correspondence (50) motivates us to introduce special configurations for the spin $S^{a}_{\text{\tiny{TD}}}$, depending on whether it is perpendicular to or aligned with the angular momentum ${\mathcal{L}}^{a}$. Taking Eqs. (14) and (50) into account and that $L^{2}=\mathfrak{Q}={\cal L}^{2}+O(\epsilon)$ with Eq. (47), we then define the following spin configurations:
@@ II Hamiltonian system on $\mathcal{N}$ and Andoyer variables
Andoyer variables, named after the French astronomer Henri Andoyer (1862-1929), are a set of canonical coordinates used in the Hamiltonian mechanics of Newtonian rigid body motion: see, e.g., Refs. [130, 131]. They are a generalization of the more well-known Hill and Delaunay variables used in the reduction of the Newtonian two-body problem [cf. Refs. [132, 91]]. The introduction of Andoyer-type variables in general relativity for spinning dynamics presented here is new, and similar to the definition of orbital elements defined in the geodesic case in Sec. II.1, although more geometry is involved. Quite remarkably, this “classical” change of coordinates can also be applied to the present relativistic context. It is, in essence, the only means by which we are practically able to transition to the physical phase space $\mathcal{P}$.
@@ II.1 Schwarzschild geodesics
The motion of a non-spinning test particle in a given background spacetime is described by geodesics. These geodesics have a well-known Hamiltonian formulation [cf. Refs. [68, 75]] on a $8$D phase space $\mathcal{G}$ [cf. FIG. 1], spanned by the $4$ spacetime coordinates $x^{\alpha}$ and their $4$ conjugated momenta $p_{\alpha}$, corresponding to the particle’s momentum $p_{a}$ components. The Hamiltonian generating the geodesic equations is simply $H_{\mathcal{G}}:=\tfrac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}$. For the case of the Schwarzschild spacetime, this reads
@@
where $(t,r,\theta,\phi)$ are the Schwarzschild-Droste coordinates and $(p_{t},p_{r},p_{\theta},p_{\phi})$ are the components of the $1$-form $p_{a}$ in the natural basis $(\partial_{\alpha})^{a}$. Just like the Kepler problem in Newtonian gravity (of which the Hamiltonian is given by the leading order in a $M/r\ll 1$ expansion of Eq.(53)), Schwarzschild geodesics are always spatially confined to a $2$-dimensional plane, reflecting the spherical symmetry. In terms of Hamiltonian mechanics, this geometric feature is nothing but a $1:1$ resonance between the $\theta$ and $\phi$ motion; this fact is not so obvious by direct inspection of Eq. (53). Because the confinement is rarely presented as a resonance in relativistic mechanics, we present a detailed discussion of it in App. A.2.
@@
where $\nu$ is the angle defining the line of nodes, where the orbital plane intersects the equatorial plane, and $\ell$ is the true anomaly, which specifies the location of the particle within the orbital plane; see Fig. 2. With the help of some Euclidean trigonometry, as detailed in App. A, we find the following formulae
@@ II.2 Construction of relativistic Andoyer variables
In the preceding section, we examined the non-spinning (geodesic) situation. In the current section, we will generalise the transformation given in Eq. (54) to the case of a spinning body, maintaining consistency at the linear order in spin. In addition to the orbital angles $\theta,\phi$, such a transformation must now involve the spin angle $\sigma$, because (the square norm of) the angular momentum $3$-vector $L^{2}$ in Eq. (42c) depends not only on $\pi_{\theta}$ and $\pi_{\phi}$, but also on the spin conjugated momentum $\pi_{\sigma}$.
This section’s aim is to achieve two main objectives: firstly, to to define three new angles $(\nu,\ell,s)$, with their conjugated momenta $(\pi_{\nu},\pi_{\ell},\pi_{s})$ related to the angular momentum of the system; and secondly, to establish a canonical transformation
$(\theta,\pi_{\theta},\phi,\pi_{\phi},\sigma,\pi_{\sigma})\mapsto(\ell,\pi_{\ell},\nu,\pi_{\nu},s,\pi_{s})\,$ (58)
between the old variables and the new coordinates. This will facilitate a natural extension of the transformation in the godesic case in Eq. (54).
In order to achieve this, it is necessary to ensure that the new conjugated momenta $(\pi_{\nu},\pi_{\ell},\pi_{s})$ are related to the angular momenta of the system by $(\pi_{\nu},\pi_{\ell},\pi_{s}):=(\pi_{\phi},L,\pi_{\sigma})$. There are alternative options that could be considered, but this particular approach provides the most straightforward form of the transformation, which bears resemblance to a well-known transformation in classical mechanics, namely the so-called Andoyer transformation [cf. Refs. [130, 131]].
@@
First, we define the set of Andoyer angles $(\nu,\ell,s)$, where $\nu$ and $\ell$ are the same angles as in the geodesic construction [cf. Sec. II.1], while $s$ belongs to the Body plane. The Andoyer angles can be related to the old ones $(\theta,\phi,\sigma)$ using spherical trigonometry identities (obtained in the spherical triangle formed by the three pairwise intersections of each plane, as in Fig. 3). These read  [^12] Compared to Ref. [131], we have shifted the angle $\ell$ by $\pi/2$ so that it coincides with its usual construction. This does not break the symplecticity because we have $\{\ell,F\}=\{\ell+\pi/2,F\}$ for any function $F$ of the phase space.
@@
The whole set of relations in Eqs. (61), (62) and (63) defines a canonical change of coordinates, expressed in the form of Eq. (58). This generalizes the geodesic transformation in Eq. (54). To prove this, it suffices to either compute the symplectic form just like Andoyer did in classical mechanics (see, e.g., Ref. [131]), or to use brute force and check that all Poisson brackets remain canonical. It is worth noting that some formulae, like Eqs. (61) or Eq. (62), only define the new Andoyer variables up to signs. These signs are all fixed by asking that the transformation be canonical. [^13] For example, only the square of $\pi_{\theta}$ can be extracted from Eq. (62) After imposing $\{\theta,\pi_{\theta}\}=1$, one readily finds $\pi_{\theta}=-\pi_{\ell}\sin\iota\cos(\nu-\phi)$. Similar features happen for other variables.
@@ II.3 Interpretation of Andoyer variables
The Andoyer variables of (non-relativistic) classical mechanics have to do with the evolution of a free rigid body [cf. Refs. [133, 130, 131]]. In our setup, they do not have quite the same Euclidean interpretation, and it is rather the mathematical formulae that they satisfy that we exploit. In this section we get a grasp at what our Andoyer-type variables represent for the relativistic system at hand.
@@ II.3.2 Planar (and equatorial) orbits
The physical significance of the angle coordinate $(s,\pi_{s})$ is most clearly elucidated by examining the limit as $\pi_{s}=0$, with the understanding that $\pi_{s}=\pi_{\sigma}$ [cf. Eq. (62)] and that this does not render all spin terms as zero [cf. Eq. (29d)]. In this limit Eqs. (61) become   [^14] In the final two equations of Eqs. (65), we also set $s=0$. This is because the dynamics will ultimately demonstrate that if $\pi_{s}=0$, then $s=0$. This result will be presented in detail in Sec. II.5
@@
$\sin\sigma\sin\theta$ $=\sin\iota\cos\ell\,,$ (65d)
$\cos\sigma\sin\theta$ $=\cos\iota\,,$ (65e)
where, in a similar manner to the inclination angle $\iota$, we assume (without loss of generality) that $\delta\in[0,\pi]$, so that $\pi_{s}=0$ is equivalent to $\cos\delta=0$ [cf. Eq. (63)], which implies both $\delta=\pi/2$ and $\sin\delta=1$.
In the first three equations, we identify the geodesic formulae in Eqs. (55), for which the motion was planar. Indeed, if $\pi_{s}=0$, then it follows from Eq. (60) that $\vec{L}$ is orthogonal to $\vec{e}_{r}$ at all times. The motion is confined into the invariant Andoyer plane that is orthogonal to $\vec{L}$. The variable $\pi_{s}$ thus controls the non-planarity of the orbit, and the Andoyer plane can be characterized as a “mean orbital plane”. In particular, the angle between the position vector $\vec{e}_{r}$ and the Andoyer plane is $\delta-\pi/2$.
The angle $s$ is measured within the Body plane, and locates the position of the spin vector $\vec{S}$ [cf Eq. (20)] with regard to the intersection of the Body and Andoyer plane. The angle $s$ coincides with the spin angle $\sigma$ only for equatorial orbits, where one has $(\theta,\phi,\iota,\sigma)=(\pi/2,\ell+\nu,0,0)$. Therefore, $s$ is merely the generalization of the angle $\sigma$, which is used to parameterise the spin vector $\vec{S}$ in the spherical triad. For further details on its physical interpretation, see Sec. III.B in Paper I for its physical interpretation.
Although the orbit of the spinning body is generally non-planar, the dynamics still exhibit an invariant plane, which is orthogonal to the constant angular momentum vector $\vec{L}$. This plane is characterized by a line of nodes $\nu$ and an inclination $\iota$ [cf. Eq. (63)]. We stress, again, that the motion of the spinning body is not confined to this plane, but stays $O(\epsilon)$-close to it.